Status Report

NOAA GOES SATOPS Morning Report 16 Jan 2001

By SpaceRef Editor
January 16, 2001
Filed under ,

SATOPS MORNING REPORT January 16, 2001

POES

NOAA-11 rev 63451 / F at 0832z on January 13: SBUV (Solar Backscatter
Ultraviolet Spectrometer) began operating in limited range at approximately
0601z.  On rev 63455 at 1540z on January 13 crew noted that SBUV grating
drive was unstuck.

NOAA-14 rev 31127 / W  at 0900z on January 13: G3A was noisy and
only 95% was recovered.  Repositioned for replay and recovered 100%
on rev 31128.

NOAA-15 rev 13879 / W at 1233z on January 13: G3B noisy from spacecraft
and 96.8% recovered. Repositioned for replay and recovered 100% on rev
13880.

NOAA-14 rev 31130 / F at 1359z on January 13: L5B noisy from spacecraft
and 92.8% was recovered.  Further recovery attempts not possible due
to spacecraft recorder management sequence constraints.

NOAA-14 rev 31131 / F at 1540z on January 13: L1B noisy from spacecraft
and 96.6% was recovered.  Further recovery attempts not possible due
to spacecraft recorder management sequence constraints.

NOAA-14 rev 31148 / W at 2013z on January 14: G5A noisy from spacecraft
and 98.1% was recovered.  No re-ingest required.

NOAA-14 rev 31158 / F at 1336z on January 15: L5B noisy from spacecraft
and 97.8% was recovered.  No re-ingest required.

NOAA-14 rev 31159 / F at 1516z on January 15: L2B noisy from spacecraft
and 97.4% was recovered.  No re-ingest required.

DMSP

OPS47 rev 29959 / FAIR at 0146z on January 13: No Thermal Fine data
to playback, not recorded due to spacecraft/ground command problems.

OPS47 rev 29960 / HULA at 0316z on January 13: Smoothed data from previous
rev recovered, but CAT1 data loss declared by user due to timelines not
being met.

OPS49 rev 05624 / POGO at 1151z on January 13:  Telemetry flagouts
are related to the recent Precision Attitude Determination Software diversion.

OPS48 rev 19522 / POGO at 1907z on January 15: Unable to link to Thule,
Greenland Remote Tracking Station (POGO).  GE Americom notified, took
pass at Fairbanks.

OPS47 rev 29998 / POGO at 2003z on January 15: Unsuccessful link with
site.  Moved pass to Fairbanks.

OPS49 rev 05658 / POGO at 2128z on January 15: Unsuccessful link with
site.  Moved pass to Fairbanks.

OPS47 rev 30004 / FAIR at 0615z on January 16: After site change due
to POGO problems, error in reading main event time resulted in short data
playback: smooth record was not going to shut off in time to recover the
data, so the recorder was shut off 2 min early.  Resulted in a Cat.
1 Data loss declared by user.

OPS47 rev 30005 / POGO at 0804z on January 16: Could not get link out
to site; no other sites visible. Cat. 1 Data loss declared by user.

OPS47 rev 30006 / POGO at 0946z on January 16: Could not get link out
to site; no other sites visible. Cat. 1 Data loss.

OPS48 rev 19531 / POGO at 1029z on January 16: Delete with no other
site available to take support.

Thule Tracking Station (POGO) comm problem summary:  GE Americom
was notified at 1920z of commanding problem with circuit from SOCC to POGO.
Job repair ticket opened. Loose cable on equipment at Thule was the problem. 
Link operational as of 1230z on January 16.

GOES

No change in the status of the GOES spacecraft.

GOES-10:  Rapid Scan Operations from 1603z on January 12 until
0003z on January 13, supporting NWS Monterey request, Southern California
severe convector weather.

At 1142z on January 13: GOES-8 image frame break due to loss of signal
synch.

At 1906z on January 13: GOES-8 North scan inhibit for lunar intrusion.

At 0320z on January 14: GOES-8 image frame break due to address phasing
error.

At 0951z on January 14: GOES-8 image frame break due to loss of signal
synch.

At 2312z on January 13: GOES-10 South scan inhibit for lunar intrusion.
Back to dual chord at 0251z on January 14 after lunar intrusion.

Questions or comments to [email protected]

SpaceRef staff editor.